OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5 ( , )

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@mathmath333

zepdrix (zepdrix):

These are lines? Just straight lines? So we can use our handy approaches to come up with equations for each line.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Given the data for the first line, \(\Large\rm (-1,9)\) and \(\Large\rm (3,-3)\), Our slope formula gives us:\[\Large\rm m=\frac{y_2-y_1}{x_2-x_1}=\frac{-3-9}{3--1}=\frac{-12}{4}=-3\]So the slope of that first line is -3.\[\Large\rm y=mx+b\]\[\Large\rm y=-3x+b\]To find the y-intercept, let's plug one of the points in and solve for b.\[\Large\rm 9=-3(-1)+b\]Solving for b gives us b=5. So the equation of our first line is, \(\Large\rm y=-3x+5\).

zepdrix (zepdrix):

Do you understand the process? :o Think you can apply that to the second set of data?

zepdrix (zepdrix):

Woops :) Not b=5, I shoulda said b=6. hehe was moving too fast.

OpenStudy (anonymous):

i know how to fid the slope but i dont understand how to find the y intercept

OpenStudy (anonymous):

find*

zepdrix (zepdrix):

What's the slope of the second line? Take a minute to calculate it :3 imma go grab a drink real quick

OpenStudy (anonymous):

the slope is 6/3 for line 2

zepdrix (zepdrix):

Mmmm ok very good.\[\Large\rm y=2x+b\]

zepdrix (zepdrix):

So we have a line,\[\Large\rm \color{orangered}{y}=2\color{royalblue}{x}+b\]And to find the y-intercept, we need to plug in a point that the line passes through. Let's use this point which was given to us: \(\Large\rm (\color{royalblue}{x},\color{orangered}{y})=(\color{royalblue}{2},\color{orangered}{5})\)

zepdrix (zepdrix):

Do you see WHERE we'll plug the values in? I tried to color code them.

OpenStudy (anonymous):

5=2(2)

OpenStudy (anonymous):

5x=2(2)

zepdrix (zepdrix):

We're not plugging anything in for the b, so don't let that part just disappear.\[\Large\rm \color{orangered}{y}=2\color{royalblue}{x}+b\]\[\Large\rm \color{orangered}{5}=2\color{royalblue}{(2)}+b\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5=4+b

zepdrix (zepdrix):

mhm, solve for b.

OpenStudy (anonymous):

but how thats what im confused about?

zepdrix (zepdrix):



zepdrix (zepdrix):

To isolate the b, we're just applying a simple algebra step. The 4 is being ADDED to the b, We want to apply the inverse operation to get the b alone. So we'll SUBTRACT 4 from each side, yah?